# A fast, flexible simulation framework for Bayesian adaptive designs - the R package BATSS

### v1

@@ Abstract
The use of Bayesian adaptive designs for randomised controlled trials has been hindered by the lack of software readily available to statisticians. We have developed a new software package (Bayesian Adaptive Trials Simulator Software - BATSS) for the statistical software R, which provides a flexible structure for the fast simulation of Bayesian adaptive designs for clinical trials. We illustrate how the BATSS package can be used to define and evaluate the operating characteristics of Bayesian adaptive designs for various different types of primary outcomes (e.g., those that follow a normal, binary, Poisson or negative binomial distribution) and can incorporate the most common types of adaptations: stopping treatments (or the entire trial) for efficacy or futility, and Bayesian response adaptive randomisation - based on user-defined adaptation rules. Other important features of this highly modular package include: the use of (Integrated Nested) Laplace approximations to compute posterior distributions, parallel processing on a computer or a cluster, customisability, adjustment for covariates and a wide range of available conditional distributions for the response.
keywords: Bayesian adaptive trial, clinical trial simulation, adaptive trial, INLA, R package, multi-arm multi-stage, Bayesian response adaptive randomisation
Dominique-Laurent Couturier1,⋆,†, Elizabeth G Ryan2,⋆, Rainer Puhr2,
Thomas Jaki1,3 and Stephane Heritier2
${}^{\star}:$ Authors contributed equally,
@@
:${}^{1}:$ MRC Biostatistics Unit, University of Cambridge, UK,
:${}^{2}:$ School of Public Health and Preventive Medicine, Monash University, Australia,
:${}^{3}:$ Faculty of Informatics and Data Science, University of Regensburg, Germany
@@ 1 Introduction
Bayesian adaptive designs are increasingly being used to deliver innovative and complex clinical trials with the aim of providing more efficiency and flexibility (e.g., Cellamare et al., 2017; Angus et al., 2020; Denholm et al., 2022; Gajewski et al., 2022). These designs combine features of Bayesian inference that are attractive for trials, such as the ability to incorporate prior information like historical controls and to express treatment effects as probability distributions, with the capability to conduct pre-planned analyses on accumulated data at scheduled time points. This allows for alterations to study design components, such as sample size, randomisation allocation ratios, and the list of available treatments in the trial, based on the results of these analyses.
@@
Multi-arm, multi-stage (MAMS) designs (Royston et al., 2003, 2011; Magirr et al., 2012; Jaki, 2015) are a type of adaptive clinical trial design that allow for multiple treatments to be compared within the same trial, often to a common control treatment, and aim to select the best treatment(s). This is achieved by performing adaptations such as: early stopping of treatment arms (removal from the study) for efficacy or lack-of-benefit (referred to here as “futility”), early stopping of the trial for efficacy or futility, and updating the randomisation probabilities based on the outcome data, to achieve certain experimental goals (known as “response adaptive randomisation”). Whilst MAMS designs can be conducted in both Bayesian and frequentist frameworks, we focus here on Bayesian MAMS designs which are naturally suited for adaptations as they can be informed by probability statements on treatment effects (which can be updated at each interim analysis), a feature not accommodated by frequentist MAMS designs. Most of the effort in terms of (open source) software development has mainly focused on the latter (e.g., nstage (Barthel et al., 2009; Blenkinsop and Choodari-Oskooei, 2019), MAMS (Jaki et al., 2019), rpact (Wassmer and Pahlke, 2023), multiarm (Grayling and Wason, 2020), nstagebin (Bratton, 2014), and DESMA (Grayling, 2019)).
Bayesian adaptive designs require extensive work when planning the trial design. Analytic sample size calculations are not feasible for such designs as closed-form solutions usually do not exist. Instead, a comprehensive set of simulations must be performed to study the operating characteristics of a proposed design (such as the type I error, power, expected sample size), the results of which are often provided in the study protocol (or a simulation appendix).
These simulations typically consider a wide range of scenarios, with varying design parameters such as the treatment effect(s), the types of adaptations present, the timing of the adaptations, the maximum sample size, or the decision rules or thresholds for making adaptations. Simulations may also be needed to fine-tune the analysis model and prior distributions, and to ensure that the design is sufficiently robust to departures from the original trial assumptions, such as the recruitment rates or outcome distribution. For each scenario, (tens to hundreds of) thousands of simulations are performed to estimate properties such as the type I error and power with sufficient precision, and to understand the potential consequences and gains of the adaptations, as well as the trial settings. Several rounds of simulations are often performed during the planning of a Bayesian adaptive trial, during which the design is refined, and further simulation work may be required during the trial. Thus, high levels of computational performance are essential to enable the software to perform these complex computations in a relatively short amount of time.
One of the main deterrents to using Bayesian adaptive designs is the paucity of easily-accessible, user-friendly software to perform the computationally-intensive design simulations. Currently, simulation of Bayesian adaptive designs generally requires a trial’s statistical design team to construct custom code or to use stand-alone simulator software (discussed below), which can be time-consuming.
@@ 1.2 Aim
The aim of this work is to provide an open-source, script-based, flexible structure (via an R software package) that can be used as a general design tool for the fast simulations of Bayesian adaptive trials. We will illustrate how our Bayesian Adaptive Trials Simulator Software (BATSS) R package can be used to define and evaluate the operating characteristics of a Bayesian adaptive trial for different types of endpoints given some common adaptations: stopping arms (or the entire trial) for efficacy or futility and response adaptive randomisation. To permit generalisability, we use a modular approach where the user can select the adaptations that they wish to include and specify their own adaptation rules through user-defined R functions. We provide several examples of such functions to assist the user.
@@ 2 Laplace approximations for the posterior distribution
For each iteration of the Monte Carlo simulations (e.g., 1…M=10’000 trials) that are performed to understand the operating characteristics of a Bayesian adaptive trial design, the posterior distribution must be computed (often on multiple occasions within a simulated trial for each interim analysis that is performed). Thus, fast computations of the posterior distribution are essential. Calculation of the posterior distribution can often involve high-dimensional integration for which closed-form expressions do not exist and so numerical techniques are needed.
Sampling-based methods, such as Markov Chain Monte Carlo (MCMC) methods, have commonly been used for estimating the posterior distribution and related quantities for Bayesian inference. These methods require large numbers of MCMC samples, which can be computationally demanding (and sometimes prohibitive), and assessment of convergence. To circumvent these problems, we use a different approach, which is deterministic and potentially much faster.
@@ 2.1 Integrated Nested Laplace Approximations (INLA)
Our package relies on the Integrated Nested Laplace Approximations (INLA) framework (Van Niekerk et al., 2023; Brown et al., 2021; Rue et al., 2009) that provides approximate analytical posteriors by assuming that the model follows a specific structure – the latent Gaussian model – and uses sparse matrices (Rue et al., 2009). Latent Gaussian models represent a wide class of commonly-used statistical models, including generalised linear (mixed) models, spatio-temporal models, survival and joint longitudinal-survival models. Details about the modelling structure and default priors can be found in (Couturier et al., 2024).
@@ 3 The BATSS R package
The aim of the BATSS R package (Couturier et al., 2024) is to provide a user-friendly tool for the fast simulation of Bayesian adaptive trial designs so that a design’s operating characteristics can be determined. It is intended that the BATSS package can be used for phase II-IV trials with interim analyses for stopping arms or the trial and alterations to the sample size or allocation ratio. As a first step, we focus our attention on MAMS designs with a common control arm (where we are primarily interested in pairwise comparisons to the control), implemented in the Bayesian framework.
@@ 3.2 Overview of the BATSS framework
[…]
@@
8:        Simulate $m$ responses [var, var.control, link]
@@
10:        Calculate target parameter posterior probabilities using the $\delta$s[which, delta.eff, delta.fut]
@@
22:        Update $m$, $n_{curr}$, and $prob$
@@
                 var.control        = list(y = list(size = 1/2)),
                                    #size = dispersion/shape parameter
@@
    summary(sim, full = TRUE)
Figure 2 shows the parameter estimates (posterior mean) against the sample size for each arm and the decision at the time of stopping the arm for the simulated design, which can be obtained using:
    plot(sim,type="estimates")
Figure 2: Scatterplot of target parameter estimates (posterior mean) versus sample size for each intervention arm (separate plots) at the time of stopping the arm, along with the decision at the time of stopping for the simulated scenario for Case Study 1 across the 10’000 simulated trials. The timing of the stop is also provided (“Early stopping” = stopping at an interim analysis).
The distribution of sample sizes for each treatment arm can also be examined using the plot function:
    plot(sim,type="size")
Figure 3: Box plot showing the distribution of overall sample size and sample size for each treatment arm across the 10’000 simulated trials for the simulated scenario for Case Study 1.

### v2

@@ Abstract
The use of Bayesian adaptive designs for randomised controlled trials has been hindered by the lack of software readily available to statisticians. We have developed a new software package (Bayesian Adaptive Trials Simulator Software - BATSS) for the statistical software R, which provides a flexible structure for the fast simulation of Bayesian adaptive designs for clinical trials. We illustrate how the BATSS package can be used to define and evaluate the operating characteristics of Bayesian adaptive designs for various different types of primary outcomes (e.g., those that follow a normal, binary, Poisson, or negative binomial distribution) and can incorporate the most common types of adaptations: stopping treatments (or the entire trial) for efficacy or futility, and Bayesian response adaptive randomisation - based on user-defined adaptation rules. Other important features of this highly modular package include: the use of (Integrated Nested) Laplace approximations to compute posterior distributions, parallel processing on a computer or a cluster, customisability, adjustment for covariates and a wide range of available conditional distributions for the response.
Dominique-Laurent Couturier1,†, Rainer Puhr2, Stephane Heritier2,
Thomas Jaki1,3 and Elizabeth G Ryan2
@@
$^{1}$: MRC Biostatistics Unit, University of Cambridge, UK,
$^{2}$: School of Public Health and Preventive Medicine, Monash University, Australia,
$^{3}$: Faculty of Informatics and Data Science, University of Regensburg, Germany
@@ 1 Introduction
Bayesian adaptive designs are increasingly being used to deliver innovative and complex clinical trials with the aim of providing more efficiency and flexibility (e.g., Cellamare et al., 2017; Angus et al., 2020; Denholm et al., 2022; Gajewski et al., 2022). These designs combine features of Bayesian inference that are attractive for trials, such as the ability to incorporate prior information like historical controls and to express treatment effects as probability distributions, with the capability to conduct pre-planned analyses on accumulated data at scheduled time points. This allows for alterations to the study design components, such as sample size, randomisation allocation ratios, and the list of treatments available in the trial, based on the results of these analyses.
@@
Multi-arm, multi-stage (MAMS) designs (Royston et al., 2003; Royston et al., 2011; Magirr et al., 2012; Jaki, 2015) are a type of adaptive clinical trial design that allow for multiple treatments to be compared within the same trial, often to a common control treatment, and aim to select the best treatment(s). This is achieved by performing adaptations such as: early stopping of treatment arms (removal from the study) for efficacy or lack-of-benefit (referred to here as “futility”), early stopping of the trial for efficacy or futility, and updating the randomisation probabilities based on the outcome data, to achieve certain experimental goals (known as “response adaptive randomisation”). Whilst MAMS designs can be conducted in both Bayesian and frequentist frameworks, we focus here on Bayesian MAMS designs which are naturally suited for adaptations as they can be informed by probability statements on treatment effects (which can be updated at each interim analysis), a feature not accommodated by frequentist MAMS designs. Most of the effort in terms of (open source) software development for MAMS designs has been done in the frequentist framework (e.g., nstage (Barthel et al., 2009; Blenkinsop and Choodari-Oskooei, 2019), MAMS (Jaki et al., 2019), rpact (Wassmer and Pahlke, 2023), multiarm (Grayling and Wason, 2020), nstagebin (Bratton, 2014; Choodari-Oskooei et al., 2023), and DESMA (Grayling, 2019)).
Bayesian MAMS and other Bayesian adaptive designs rely on Bayesian inference to adapt trials. At each interim analysis, Bayesian methods combine prior information on model parameters (which might be informed by clinicians, past information, or be vague or non-informative) and the likelihood of parameters given the observed data, resulting in a posterior multivariate probability distribution that summarises the known information about the parameters at a given stage. This distribution allows decision making (such as determining trial or arm successes) to be based on parameter-specific probability statements such as ‘there is a 95% probability that the new treatment is superior to the standard treatment’.
1.1 Monte Carlo Simulations
Whilst traditional designs with a single endpoint may rely on (asymptotic) statistical theory to ensure suitable power and control of type I error, such analytical results are typically not available for complex adaptive designs. Consequently, trial operating characteristics (such as power, type I error, and expected sample size) are typically estimated using Monte Carlo simulations in such cases. Among the many research papers and guidelines describing how to perform and evaluate simulation studies (see, e.g., Morris et al. (2019); White et al. (2023); Ankolekar et al. (2020); Yuxin Ding and Qu (2025)), we focus here on those by
  • the ‘International Council for Harmonisation of Technical Requirements for Pharmaceuticals for Human Use’ (ICH), especially ‘ICH-E20’ the new draft guidelines dedicated to adaptive designs $I_{1}$, (ICH 2025)] and ‘ICH-E9R1’ the draft of the revision of the guidelines dedicated to multiplicity corrections $I_{2}$, (ICH, 2025)],
  • the ‘European Medicines Agency’ (EMA), especially the agency initial reflections on adaptive designs $E_{1}$, (EMA 2007)] and questions and answers dedicated to complex clinical trials $E_{2}$, (EMA, 2022)],
  • the US ‘Food and Drug Administration’ (FDA), especially the agency initial reflections on adaptive Design $F_{1}$, (FDA 2019)] and questions and answers dedicated to complex clinical trials $F_{2}$, (FDA, 2020)],
as they provide detailed guidance on expectations to ensure regulatory acceptance and uptake of complex clinical trials, particularly Bayesian ones. For simplicity, to cite these documents, we will use $X^{p}_{d}$ to refer to page ‘p’ (subscript) of the document ‘d’ (lower script) of source ‘X’, where ‘X’ can either be ‘E’, ‘F’ or ‘I’, respectively corresponding to the EMA, FDA and ICH in the remaining part of this Section. The aforementioned guidelines all mention the need for extensive computer simulations to determine important aspects of the design on which the assessment of design proposals is to be made. Such simulations should be “carefully planned, conducted and reported” ($I_{1}^{5}$) and include, notably,
  • the careful evaluation of the “chance of producing wrong conclusions” ($F^{4}_{1}$) where the probability of erroneous conclusions should be assessed by estimating the frequentist operating characteristics of type I error by computer-intensive simulations ($F^{6}_{1}$, $E^{6}_{1}$). When multiple null hypotheses are being tested, the probability of rejecting at least one of them is often controlled ($F^{19}_{2}$, $E^{5}_{2}$), sometimes referred to as the family-wise error rate (FWER). The EMA considers type I error control to be “a minimum prerequisite for statistical methods to be accepted in the regulatory setting” ($E^{5}_{1}$),
  • a detailed study of other operating characteristics like power and expected, minimum and maximum sample sizes, for example ($F^{11}_{2}$$I^{21}_{1}$),
  • the optimisation, typically by simulation, of the timing and number of interim analyses ($F^{6}_{1}$, $F^{18}_{2}$),
  • a comparison of the above operating characteristics with the ones obtained when considering alternative designs, including conventional (i.e., non-adaptive) ones ($F^{3}_{1}$, $I^{20}_{1}$).
  • sensitivity analyses to different chosen aspects of the design to ensure operating characteristics are robust to deviations of the assumptions made, notably,
    • prior choices ($F^{4}_{1}$, $I^{25}_{1}$), especially when estimating the (frequentist-like) type I error and when considering informative priors via the borrowing of external information (past data, expert opinion). In that case, “it is also important to evaluate the current trial data with no borrowing” ($I^{26}_{1}$).
    • the anticipated timing of interim analyses ($I^{11}_{1}$), related to, for example, changes to Data Monitoring Committee meeting dates,
    • nuisance parameter values ($F^{18}_{2}$, $I^{21}_{1}$), for example,
        the assumed variance for a Gaussian primary endpoint given the predictors (treatment effects and confounders),
        the assumed shape parameter when modelling overdispersed counts with a negative binomial distribution, as well as the average count number in the control group when targeting a given treatment-related fold change,
        the probability of success in the control group when modelling a binary endpoint by means of a binomial distribution and targeting a given treatment-related odds ratio as effect size,
    • drop-out and missingness levels ($I^{20}_{1}$),
    • the use of non-binding futility stopping rules when binding ones were planned, as data monitoring committees sometimes opt for flexibility and might consider the futility stopping criteria as a “guideline that can be deviated from” ($I^{12}_{1}$). Of note, in such cases, the FDA would likely consider the trial “to have failed to provide evidence of efficacy, regardless of the outcome of the final analysis” ($F^{12}_{2}$), hence the importance of such sensitivity analyses.
Other relevant aspects of these guidelines include
  • the suggestion, to mitigate the fact that an infinite number of scenarios are compatible with the selected null and alternative hypotheses ($F^{18}_{2}$, $I^{22}_{1}$), to determine a limited set of scenarios that adequately represent plausible scenarios ($F^{18}_{2}$, $I^{13}_{1}$, $I^{22}_{1}$). This could be achieved, for example, by relying on medical expert opinion ($F^{19}_{2}$). The aim would then be to make sure that the maximum type I error across the grid of scenarios does not exceed the target error rate, for example, 2.5% for a one-sided test ($F^{6}_{1}$), and that the power exceeds the target one in all selected cases.
  • details on the suitable number of Monte Carlo trials per scenario and the way to select seeds per scenario: To achieve a suitable level of precision, it is noted that 100,000 Monte Carlo trials would be sufficient “in most cases” when evaluating the type I error, and that 10,000 simulated trials might be enough to evaluate power ($F^{19}_{2}$, $I^{22}_{1}$). The use of different seeds per scenario is also recommended “to avoid [observing] consistently atypical results across scenarios” ($F^{19}_{2}$).
@@ 1.3 Aim
The aim of this work is to provide an open-source, script-based, flexible structure (via an R software package) that can be used as a general design tool for the fast simulations of Bayesian adaptive trials and that would enable the efficient evaluation of design characteristics across a wide range of scenarios and number of simulations, as required by the EMA, FDA, and ICH guidelines mentioned in Section 1.1.
We will illustrate how our Bayesian Adaptive Trials Simulator Software (BATSS) R package can be used to define and evaluate the operating characteristics of a Bayesian adaptive trial for different types of endpoints given some common adaptations: stopping arms (or the entire trial) for efficacy or futility and response adaptive randomisation. To permit generalisability, we use a modular approach where the user can select the adaptations that they wish to include and specify their own adaptation rules through user-defined R functions. We provide several examples of such functions to assist the user.
@@ 2 Laplace approximations for the posterior distribution
In the Bayesian generalised linear models (GLM) considered in this work, the response vector of $n$ observations, $\mathbf{y}=[y_{1},...,y_{n}]^{T}$, is assumed to follow a density function depending on parameters $\bm{\mu}$} and $\phi$
$\mathbf{y}\sim f(\mathbf{y}\ |\ {\bm{\mu}},\ \phi)$
where $f(\cdot)$ is a density distribution belonging to the exponential family (or similar), $\bm{\mu}$} denotes the expectation of $\mathbf{y}$, and $\phi$ is an optional nuisance parameter. The exponential family encompasses a wide range of distributions like the binomial distribution for binary endpoints, the Poisson and negative binomial distribution for count data, and the beta, Gaussian and gamma distributions for continuous endpoints. In GLMs, the response mean is linked to the model predictors through a link function:
${\bm{\mu}}=g(\mathbf{X}{\bm{\beta}})$
where $\mathbf{X}$ denotes the ($n\times p$) predictor design matrix, and $\bm{\beta}$} denotes the regression parameter vector whose interpretation depends on the chosen link function. The expit function ${\bm{\mu}}=\exp(-\mathbf{X}{\bm{\beta}})^{-1}$ is often used for endpoints bounded between 0 and 1 (like binary or beta endpoints), leading to $\bm{\beta}$} coefficients corresponding to odd ratios. The exponential function ${\bm{\mu}}=\exp(\mathbf{X}{\bm{\beta}})$ is typically used for strictly positive endpoints (like counts), leading to $\bm{\beta}$} coefficients corresponding to fold changes. The linear model is a special case of GLM in which $f(\cdot)$ corresponds to the Gaussian distribution and ${\bm{\mu}}=\mathbf{X}{\bm{\beta}}$ (identity link function).
In the frequentist framework, the GLM parameter vector combining the $\bm{\beta}$} regression parameters and optional nuisance parameter $\phi$, ${\bm{\theta}}=[{\bm{\beta}},\phi]^{T}$, is estimated by maximising the likelihood function of the parameters given the data. In the Bayesian framework, the parameter vector is assumed to follow a (multivariate) prior distribution ${\bm{\theta}}\sim\pi({\bm{\theta}})$. Bayesian inference then asserts that the posterior probability distribution of the fitted model parameters is proportional to the likelihood function multiplied by the prior distribution of the parameters.
$\pi({\bm{\theta}}\ |\ \mathbf{y},\mathbf{X})\propto f(\mathbf{y}\ |\ {\bm{\theta}},\mathbf{X})\ \pi({\bm{\theta}})$
For each iteration of the Monte Carlo simulations (e.g., 1…M=10,000 trials) that are performed to understand the operating characteristics of a Bayesian adaptive trial design, the posterior distribution must be computed (often on multiple occasions within a simulated trial for each interim analysis that is performed). Thus, fast computations of the posterior distribution are essential. Calculation of the posterior distribution can often involve high-dimensional integration for which closed-form expressions do not exist and so numerical techniques are needed.
Sampling-based methods, such as Markov Chain Monte Carlo (MCMC) methods, have commonly been used for estimating the posterior distribution and related quantities for Bayesian inference. These methods require large numbers of MCMC samples which can be computationally demanding (and sometimes prohibitive), and necessitate both the assessment of convergence of the underlying Markov chain to stationarity as well as the assessment of convergence of Monte Carlo estimators to population quantities (Roy, 2020; Robert and Casella, 2010). To circumvent these problems, we use a different approach, which is deterministic and potentially much faster.
@@ 2.1 Integrated Nested Laplace Approximations (INLA)
Laplace approximation provides an analytical method to estimate the posterior distribution of parameters in complex models by assuming that the posterior distributions can be approximated by Gaussian densities centered at their mode (or maximum). The calculation of integrals related to the posteriors are therefore greatly simplified, typically leading to computational gains, especially for complex models. R has several implementations of Laplace approximations for Bayesian inference, such as in the TMB (Template Model Builder), LaplacesDemon and mgcv packages, for example.
BATSS relies on the Integrated Nested Laplace Approximations (INLA) framework (Van Niekerk et al., 2023; Brown et al., 2021; Rue et al., 2009) that provides approximate analytical posteriors by assuming that the model follows a specific structure – the latent Gaussian model – and uses sparse matrices (Rue et al., 2009). Latent Gaussian models represent a wide class of commonly-used statistical models, including generalised linear (mixed) models, spatio-temporal models, survival and joint longitudinal-survival models. Details about the modelling structure and default priors can be found in (Couturier et al., 2025).
@@ 3 The BATSS R package
The aim of the BATSS R package (Couturier et al., 2025) is to provide a user-friendly tool for the fast simulation of Bayesian adaptive trial designs so that a design’s operating characteristics can be determined. It is intended that the BATSS package can be used for phase II-IV trials with interim analyses for stopping arms or the trial and alterations to the sample size or allocation ratio. As a first step, we focus our attention on MAMS designs with a common control arm (where we are primarily interested in pairwise comparisons to the control), implemented in the Bayesian framework.
@@ 3.2 Overview of the BATSS framework
4:  Set $prob=\color[rgb]{0,0,1}[\texttt{prob0}]$ # initial randomisation probabilities
@@
8:   Simulate $m$ responses [model, var, var.control, link]
@@
[…]
@@
22:   Update $m$, $n_{curr}$, and probprob
@@
var.control = list(y = list(size = 1/2)), # shape
@@
summary(sim, extended = TRUE)
Figure 2 shows the probability of stopping the trial at each look under the assumed parameter values as well as the expected trial sample size. Figure 3 displays the sample size of each simulated trial per group with, for the target parameters, points colour-coded and shaped according to the decision and timing when the trial stopped. Figure 4 shows the parameter estimates (posterior mean) against the sample size for each arm and the decision at the time of stopping the arm for the simulated design. These plots can be obtained using:
plot(sim, which = 2:4)
where which=1 produces the boxplot of the total and per group sample sizes of the simulated trials.
Figure 2: Barplot of the probability of stopping (y-axis) at each look (x-axis) as observed in the simulated samples. The vertical blue dashed line shows the estimate of the expected sample size (ESS).
Figure 3: Violin plot (and boxplot) of the simulated sample sizes per treatment arm for each trial and corresponding expected sample sizes. For target parameters, these points are colour-coded and shaped by the decision and timing of when the arm stopped.
Figure 4: Scatterplot of target parameter estimates (posterior mean) versus sample size for each intervention arm (separate plots) at the time of stopping the arm, along with the decision at the time of stopping for the simulated scenario for Case Study 1 across the 10,000 simulated trials. The timing of the stop is also provided (“Early stopping” = stopping at an interim analysis).The horizontal blue lines show the assumed target parameter values.